OpenStudy (anonymous):

The equation y = 90x shows that a cheetah is running at 90 feet per second. Complete the table.

OpenStudy (anonymous):

So if x = 1 you get y = 90 * 1. Put that in for the first value. What do you get if x = 2?(remeber in y = 90x you will just put a 2 where the x is)

OpenStudy (anonymous):

thanks ^.^

OpenStudy (anonymous):

Well it is important to understand what x and y actually are in this equation. And not just simply plug in a value for x, because you were told. When we are told that a cheetah runs 90 feet per second. We are given the equation Y=90X If you look at the equation you can see that for every second the cheetah will go a total of 90 feet. So if he ran for 1 second, he will have gone 90 feet. If he runs for 2 seconds he will have gone 180 feet. In this equation the variable X represents how many seconds the cheetah has run, where Y is how many feet the cheetah ran. So when x=1 Y=90(1), x=2 Y=90(2), x=4 Y=90(4) so the answer of Y will be the feet the cheetah traveled.